OpenStudy (anonymous):

A large dump truck will empty its full load of stone onto a small plot of land for storage. The rock forms a cone that is 9 feet tall. The truck holds 144 cubic feet of stone. What is the area of the circular base of the pile of stones?

OpenStudy (anonymous):

@ganeshie8 or someone please help me with this!

OpenStudy (anonymous):

answer choices.. 48 square feet.. 48/pi ... 16... or sqrt 48/pi

OpenStudy (bibby):

the full load is 144 cubic feet the formula for th e volume of a cone is \(\large \frac{1}{3}\pi r^2 h\) set up an equation knowing that the full load \((144ft^3)\) is equal to the volume of the cone where h=9. then recall that the area of a circle is \(\pi r^2\)